A sphere has a potential difference between it and the earth of 5 V when charged with 14 nanocoulomb. What is the capacitance?


1
Expert's answer
2022-03-11T08:28:39-0500

The capacitance is given as follows (see https://physics.nfshost.com/textbook/09-Capacitance/01-Definition.php):


"C = \dfrac{Q}{\Delta V}"

where "Q = 14\times 10^{-9}C, \Delta V = 5V". Thus, obtain:


"C = \dfrac{14\times 10^{-9}C}{5V} = 2.8\times 10^{-9}F"

Answer. "2.8\times 10^{-9}F".